OpenStudy (anonymous):

Turn 49/50 into a percent

OpenStudy (tyteen4a03):

49/50 = 98/100

OpenStudy (anonymous):

98%

OpenStudy (unklerhaukus):

\[\frac{49}{50}=\frac{2\times 49}{2\times 50}=\frac{98}{100}=98\%\]

OpenStudy (anonymous):

49/50=x/100 x=(49/50)*100

OpenStudy (anonymous):

the formula is answer=(x\T)*100 whwre x is the total under consideration and T is the overall total